OpenStudy (anonymous):

can someone explain y=1/2^1 -3

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 2^{1} }-3\]

OpenStudy (anonymous):

what don't you understand?

OpenStudy (anonymous):

ya I know I just don't know how to solve for y like after that what do you do

OpenStudy (anonymous):



OpenStudy (anonymous):

notice there is no brackets, so what is the 1st thing you need to do?

OpenStudy (anonymous):

multiply/divide

OpenStudy (anonymous):

I'm horrible at math so ya

OpenStudy (anonymous):

yes and also 1 thing to remember power 1st

OpenStudy (anonymous):

oh ya

OpenStudy (anonymous):

so then 2^1=2 so it would be 1/2

OpenStudy (anonymous):

I believe you can do it its simple really :)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

then \[\frac{ 1 }{ 2 }-\frac{ 3 }{ 1 }=\frac{ 5 }{ 2 } ??\]

OpenStudy (anonymous):

im horrible at this :(

OpenStudy (anonymous):